OpenStudy (2yoututu11):

The length of one leg of a right triangle is 24 ft, and the length of the hypotenuse is 40 ft. What is the length of the other leg? A. 6.3 feet B. 32.0 feet C. 46.7 feet D. 64.0 feet

OpenStudy (anonymous):

your job is to compute \[\sqrt{40^2-24^2}\]

OpenStudy (anonymous):

because you know \[a^2+b^2=c^2\] \[24^2+b^2=40^2\] \[b^2=40^2-24^2\] \[b=\sqrt{40^2-24^2}\]